# Ptychoformer: A Physics-Guided Deep Learning Framework for Ptychographic Imaging

### v1

@@ IV-E2 Performance Across Overlap Ratios
IV-F Ablation Study
To justify each component of the proposed Ptychoformer, we conducted comprehensive ablation studies. We compared several variants of our model: 1) Ours without NonLocal Coherence Branch (Ours w/o NLB), using only the Local Dependencies Branch; 2) Ours with Multi-Head Self-Attention (MHSA) in NLB (Ours w/ MinN), replacing PCA with standard MHSA in the NonLocal Branch; 3) Ours without Decoder (Ours w/o D), using fully connected layers instead of CNN for upsampling in the decoder; 4) Ours without Output Constraints (Ours w/o OC), removing the amplitude and phase constraints in the output layer; and 5) Ours (Full Model), the complete proposed Ptychoformer model. Table III presents the quantitative results of these variants, with the best performance in bold. The NonLocal Coherence Branch, reflecting diffraction patterns’ global nature, showed a modest impact, primarily affecting structural integrity. PCA demonstrated clear advantages over standard MHSA, with its replacement by MHSA causing a 4.16% drop in PSNRamp and 3.56% in SSIMamp.
The CNN decoder proved crucial, its absence leading to a substantial 23.58% decrease in PSNRamp and 30.46% in SSIMamp. This significant impact is attributed not only to its role in maintaining spatial coherence during upsampling but also to its inductive bias, which enables effective learning from limited data. Output constraints, while showing a modest quantitative impact, play a vital role in enforcing physical consistency. These results validate the effectiveness of each proposed component in our architecture.
@@ V Discussion
V-A Frequency Analysis
Further analysis of the noise-free full-stitched field using 1D diagonal cross-sections of the average 2D power spectral density (PSD) and frequency distribution statistics (Fig. 9). In Fig. 9(a), Ptychoformer’s PSD curve closely follows the ground truth, while the DPI model exhibits significant mid and high-frequency fluctuations, indicating potential artifacts. Quantitatively, Ptychoformer achieves a frequency distribution (99.75%, 0.22%, 0.03% for low, mid, and high frequencies, respectively), closely approximating the ground truth (99.58%, 0.38%, 0.04%). This performance exceeds that of CNN-based models, which show mid and high frequency degradation (e.g., PtychoNN: 99.93%, 0.06%, 0.01%). These results highlight Ptychoformer’s efficacy in extracting mid-to-high frequency information and retrieving corresponding detailed features, leading to superior artifact suppression.
Figure 9: Frequency analysis comparison of full-scene ptychographic retrievals. (a) 1D diagonal cross-sections of average 2D PSD from stitched simulations. Red curve: ground truth; Blue curve: models. The red box represents obvious abnormal model retrieval. (b) Quantitative breakdown of energy distribution across low, mid, and high frequency bands. Frequency ranges are defined based on the radial distance from the PSD center, with boundaries at 1/3 and 2/3 of the maximum frequency.
Figure 10: Performance analysis of weighted loss function $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ in ptychographic retrieval. PSNR, SSIM are plotted against $\alpha$ for amplitude and phase retrievals.
@@ V-B Loss Function Analysis
V-C Model Variants
To validate the effectiveness of our physics-informed design, we compared Ptychoformer with two state-of-the-art hybrid CNN-ViT models, adapted for ptychography task. The first is a TransUNet-inspired model [37], which uses a hybrid CNN-Transformer encoder where CNN extracts features and the Transformer processes them further. Its decoder employs a cascading structure with skip connections. The second is a SegFormer-based model [38], which features a hierarchical structure with progressive resolution reduction for multi-scale features. This model replaces the first linear layer in the feed-forward network with a $3\times 3$ convolution. We conducted the same experiment as in Section IV-B2 under noise-free conditions. Table V shows that Ptychoformer outperforms both models. TransUNet and SegFormer’s limitations stem from designs optimized for real-space image processing. The TransUNet variant, while benefiting from the combination of CNN and Transformer, still relies on local feature extraction in its initial stages. This approach is suboptimal for capturing the global coherence information present in diffraction patterns. Similarly, the SegFormer variant’s hierarchical structure, while effective for multi-scale feature extraction in natural images, may lose critical high-frequency information in the context of diffraction patterns due to its progressive downsampling.
TABLE V: Performance comparison of Model Variants.
Method Amplitude Phase
PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
TransUNet[37] 13.57 81.20 18.77 67.80
SegFormer[38] 13.76 83.30 19.12 68.10
Ptychoformer 14.42 87.00 19.41 70.60
@@ VI Conclusion
In this study, we introduced Ptychoformer, a deep learning framework designed for ptychographic retrieval from frequency-domain data. By incorporating physical insights into the model architecture, particularly through the PCA mechanism and dual-branch structure, Ptychoformer demonstrated superior performance over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency. The principles underlying Ptychoformer show promise for extension to other frequency-domain inverse imaging problems, potentially advancing high-resolution retrieval techniques across various scientific disciplines.

### v2

@@ IV-F Ablation Study
To justify each component of the proposed PPN, we conducted comprehensive ablation studies. As shown in Fig. 9, we compared both amplitude and phase reconstructions, including 2D reconstructed images and 1D cross-sectional profiles. We examined several model variants: 1) Ours without NonLocal Coherence Branch (Ours w/o NLB), which means using only the Local Dependencies Branch and removing the NonLocal Coherence Branch that contains proposed PoCA; 2) Ours with Multi-Head Self-Attention (MHSA) in NLB (Ours w/ MinN), replacing PoCA with standard MHSA in the NonLocal Branch; 3) Ours without Decoder (Ours w/o D), using fully connected layers instead of CNN for upsampling in the decoder; and 4) Ours (Full Model), the full proposed model. The ablation experiments were conducted using the same training set, test set, and parameter settings as in Section IV-B.
The quantitative results are presented in Table II. The removal of the NonLocal Coherence Branch leads to decreased structural coherence in phase reconstruction, which is clearly observable in the 1D profiles. Replacing PoCA with standard MHSA results in a 4.16% drop in PSNRamp and 3.56% in SSIMamp, with poorer edge transitions visible in the profile plots. The absence of the CNN decoder causes the most significant performance degradation, with PSNRamp and SSIMamp decreasing by 23.58% and 30.46%, respectively. The reconstructed images without the CNN decoder appear blurry, consistent with the experimental results in [17]. The 1D profiles reveal that without the CNN decoder, the model struggles to maintain accurate amplitude levels and phase transitions. The full model achieves the best performance across all metrics, particularly in regions with dramatic phase changes in the 1D cross-sectional profiles. These results validate the effectiveness of each proposed component in our architecture.
IV-G Model Variants
To validate our model’s effectiveness, we benchmarked against two widely-adopted hybrid CNN-ViT architectures from computer vision: TransUNet [34] and SegFormer [35]. These baselines remain relevant in recent research across multiple domains [36, 37, 38, 39], demonstrating their continued utility for important results in various fields.
The first variant adopts a TransUNet-inspired architecture [34], employing a hybrid CNN-Transformer encoder where CNNs extract initial features before Transformer processing, with a decoder utilizing a cascading structure and skip connections. The second implements a SegFormer-based approach [35], featuring a hierarchical structure with progressive resolution reduction for multi-scale feature extraction and replacing the first linear layer in the feed-forward network with a $3\times 3$ convolution, similar to the approach used in PtychoFormer[26]. PtychoFormer uses an encoder based on SegFormer and a decoder adapted specifically for ptychography tasks. Experiments conducted under noise-free conditions (see Section IV-B2) demonstrate that PPN outperforms all baseline models (Table III). The PtychoFormer results are comparable to our SegFormer-inspired model due to their similar structures. These variants’ limitations stem from designs optimized for real-space image processing: The TransUNet variant, while benefiting from the combination of CNN and Transformer, still relies on local feature extraction in its initial stages. This approach is suboptimal for capturing the global coherence information present in diffraction patterns. Similarly, the SegFormer variant’s hierarchical structure, while effective for multi-scale feature extraction in natural images, may lose critical high-frequency information in the context of diffraction patterns due to its progressive downsampling.
TABLE III: Performance comparison of Model Variants.
Method Amplitude Phase
PSNR (dB) SSIM (%) PSNR (dB) SSIM (%)
TransUNet[34] 13.57 81.20 18.77 67.80
SegFormer[35] 13.76 83.30 19.12 68.10
PtychoFormer[26] 13.77 83.54 19.17 68.18
PPN (Ours) 14.42 87.00 19.41 70.60
@@ V Discussion
Figure 10: Comparison of different models in terms of their complexity (parameters and FLOPS) (left) and training efficiency (convergence epochs and time per epoch) (right).
V-A Computational Complexity Analysis
Our analysis compares PPN with CNN-based methods (PtychoNN, PtyNet, DPI) and the ViT-CNN hybrid baseline (PtychoFormer[26]) through computational metrics in Fig. 10. While maintaining hierarchical CNN decoders common to all compared methods, our encoder design achieves substantial efficiency gains.
Among CNN approaches, PPN demonstrates intermediate complexity (6.12 MB parameters) between PtyNet (1.24 MB) and DPI (57.14 MB), yet converges faster (30 epochs) than both PtychoNN (45 epochs) and DPI (35 epochs). The 0.382 GFLOPs cost represents a 63.7% reduction from DPI’s 1.049 GFLOPs, confirming efficient feature learning despite transformer integration. Compared to PtychoFormer’s ViT-CNN architecture, our model reduces parameters by 91.1% (6.12 MB vs. 68.96 MB) and FLOPs by 65.4% (0.382 G vs. 1.105 G) while matching convergence speed (30 epochs). This efficiency comes from three optimizations: 1) fixed feature dimensions preventing channel inflation, 2) parallel dual-branch processing instead of hierarchical refinement, and 3) physics-inspired attention limiting learnable parameters. These changes eliminate redundant operations while preserving physical priors.
Figure 11: Performance analysis of weighted loss function $L_{\text{combined}}(\alpha)=\alpha L_{\text{MSE}}+(1-\alpha)L_{\text{NSSIM}}$ on simulated data. PSNR, SSIM are plotted against $\alpha$ for amplitude and phase retrievals.
@@ VI Conclusion
We propose PPN, a physics-inspired deep learning framework that significantly improves ptychographic reconstruction through two core innovations: (1) an architecture combining local feature extraction with global pattern coherence modeling, and (2) a novel attention mechanism tailored for diffraction physics. PPN demonstrated superior performance over existing state-of-the-art methods, particularly in high-frequency artifact suppression and data efficiency. Current limitations include performance validation primarily on simulated and small-scale experimental data – future work will focus on large-scale real-world deployments across diverse imaging conditions. The framework’s physics-inspired design principles show promising extensibility to other frequency-domain inverse problems like cryo-EM and astronomical imaging.